# MultiHal: Multilingual Dataset for Knowledge-Graph Grounded Evaluation of LLM Hallucinations

### v1

@@ 2.4 Multilinguality
For enabling multilinguality for MultiHal, we employ the Nllb-200 3.3bn costa2022no model and focus on its five well-performing European languages, namely German, Italian, French, Portuguese and Spanish. Our generation hyperparameters are specified in Table 2. Empirically, we found these hyperparameters to work the most optimal for our use case. We also noted that by separating the labels with semicolons yielded more accurate translations than having KG path labels purely whitespace separated, we attribute this to the improper grammatical structures that occur when label entities are not separated. We observe that Nllb-200’s output translations are generally of high quality, yet Nllb-200’s model does not always correctly output semicolon separation between the entities and predicates with respect to the English source.
@@ 3 Experimental Setup
The baseline experiments are set up using a simple prompt-based knowledge injection method. The prompt $\mathcal{P}$ is formatted as $\mathcal{P}=(\mathcal{K},\mathcal{Q})$, where $\mathcal{K}$ is knowledge in the form of a KG path and $\mathcal{Q}$ is the question of the data point, see Appendix F Listing 10 for KG-RAG and Listing 11 for vanilla QA for used prompts.
We conduct experiments with and without knowledge $\mathcal{K}$ (vanilla QA and KG-RAG respectively) to measure the effectiveness of the factual information contained in the KG paths. We measure the semantic similarity between ground-truths and model predictions using Multilingual-MiniLM-L12-v2666https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 wang2024multilingual , the choice of the sentence embedding model is based on results in the MMTE benchmark mmte_benchmark , its multilingual capabilities, and comparatively small parameter count. Semantic similarity is computed by mean-pooling the last hidden states per token for each sentence, applying L2 normalization and computing the dot-product between the ground-truth and LLM prediction representations. For experimental conditions, we use Gemini 2.0 Flash, GPT-4o Mini, and Llama 3.3 70bn instruct models.
@@ 4.1 Preliminary Baselines
Considering that our methodology primarily relies on an LLM-judge for filtering and rating the quality of KG paths, we conduct a preliminary baseline test for the English subset to observe the performance of LLM judges. We conduct the experiment using a subset of MultiHal, see Figure 2 for the data distribution. The goal of this is to gain insights on the expected quality of the KG paths.
@@
From the results in Table 3 we see that paths judged by GPT 4o-Mini have a higher correlation with the semantic similarity. Assuming that low quality paths (rated 1-3) would impair the LLM output quality and high quality paths (4-5) would improve it, we chose to run the full baseline experiments with GPT 4o-Mini. For a more in-depth breakdown of GPT 4o-Mini performance per dataset and domain, please refer to Appendix G.
@@ 4.2 Baseline Experiments
We report our baseline experiment results in Figure 3. The results show the mean semantic similarity along with standard deviation across the data points for a given condition. For a full numerical overview, please refer to Appendix H Table 8. The results showcase a consistent improvement of KG-RAG setting over QA, indicating that the mined KG paths are meaningful for a model to generate a more semantically correct result than without. Given that Figure 3 showcases mean score improvements in the range of 0.12 - 0.36 points with comparable standard deviations, we assume the QA and KG-RAG results as seperate distributions, therefore we do not compute formal statistical significance tests. Table 4 presents more fine-grained results per domain and dataset for each test condition, with mean performance over all languages.
@@
Gemini 2.0 Flash Llama 3.3 70b instruct GPT 4o Mini
felm (wk) 73 17 0.63 0.794 -0.164 0.742 0.789 -0.047 0.853 0.826 0.027
@@ 5 Discussion
Overall, the results from Figure 3 depict a consistent improvement for all tested LLMs in QA and KG-RAG settings; for detailed results refer to Table 8 in Appendix H. Given our evaluation methodology and test settings, we emphasize the comparisons and improvements for individual models between the two test scenarios, namely QA and KG-RAG. Therefore, we do not compare results across the models primarily due to varied parameter counts, and closed-source development of Gemini and GPT models.
@@
While scoping into specific domains in Table 4, we see that the performance fluctuates, although the foundational benchmarks SimpleQA, HaluEval, Defan and Shroom2024 contain approximately 95% of all data points for which we see consistent improvements on a per-model basis. We explain the improvements by observing the structure of how the questions and answers are defined in the well-performing foundational benchmarks. In a general case, as the Table 4 depicts the best-performing subsets, such as Defan, SimpleQA, HaluEval and Shroom2024, define the question explicitly and unambiguously with a single entity answer. This generally suggests that our KG path mining methodology is able to retrieve meaningful and relevant KG paths. Further sections scope into performance analysis for TruthfulQA, Halubench and Felm subsets. We supply some example problematic data points in Appendix I.
Temporal, Leading, Suggestive and Reasoning Questions. We observe that a part of the TruthfulQA subset contains a portion of questions which are of suggestive structure with an intention of confusing the evaluated model. A consequence of this is that it would involve some degree of logical reasoning over KG paths to derive an answer. Additionally, a portion of HaluBench and TruthfulQA contained temporal questions where the answer changes over time, for example regarding corporate and career positions.
@@
Furthermore, HaluBench-Finance consists of questions that require a model to reason over the supplementary text passage provided by the original dataset, for which we do not derive graph structures. Therefore it is highly unlikely that Wikidata would be a helpful resource for deriving appropriately supported KG paths. Our evaluation pipeline could benefit from integration of a reasoning methodology similar as per Generate-on-Graph genOnGraph or Think-on-Graph sun2023think . Refer to Appendix I Listings 15 and 13 for explicit examples.
@@
Domains. Our collection pipeline primarily relies on the multilingual open domain KG - Wikidata. For domains such as HaluBench-Pubmed, TruthfulQA-Health and HaluBench-Covid, performance can be improved by utilizing medical domain knowledge graphs, for example PubMed xu2020building or PrimeKG chandak2023building . We outline that the modularity of our pipeline allows for easy substitute of the KG endpoint.
@@
Sentence Embedding Limitations. We also note that our sentence embedding evaluation may not always accurately capture the semantics with respect to the question. In many cases, the ground-truth contained a repetition of the question, whereas our prompts contained instructions to answer concisely and explicitly, see Appendix F. Consequentially, some data points were evaluated with a relatively low semantic score even though the model responses directly, or with minimal deviations, answered the question. We note that TruthfulQA and Felm have been particularly affected by this as their ground-truth answers contain repetitions of the text but our model responses are more focused on single, explicit entities without linguistic overheads. Refer to Appendix I Listings 12 and 14
@@ 6 Related Work: Use of KGs in Datasets and Language Modeling
Multiple surveys discuss KG usage in the context of LLMs, particularly outlining future work roadmaps and synergy russaKgSurvey ; pan2024unifying ; kau2024combining , discussing KGs in context of factuality, hallucination mitigation, multilinguality lavrinovics , and graph-retrieval augmented generation peng2024graph . We identify these as useful starting points for researchers new to the topic.
Language Modeling. Previous works make use of KG structures sansford2024grapheval ; rashad-etal-2024-factalign as part of their hallucination detection methodology by extracting graph structures from a given piece of text passage. Another approach involves reasoning over KGs sun2023think or generating SPARQL queries from natural text srivastava2024mst5 . FactKG kim-etal-2023-factkg and Fleek fatahi-bayat-etal-2023-fleek propose methodologies using KGs to aid fact-checking.
All the aforementioned language modeling approaches present KG information as in-context knowledge. However, in-context knowledge has limitations — particularly when there are conflicts between LLM’s internal knowledge and the provided context, or when there is limited transparency into how the model integrates and utilizes the external knowledge. An alternative approach is to encode the information as part of the model’s weights using adapter networks pfeiffer2023modular ; tian-etal-2024-kg ; ribeiro-etal-2022-factgraph .
Factually Oriented and KG-QA Datasets. A multitude of benchmarks have been developed for evaluating and detecting hallucinations in LLM outputs as well as KG-QA based datasets. Benchmarks such as Shroom2025 shroom2025 , Felm zhao2023felm , TruthfulQA lin2021truthfulqa , simpleQA , HaluBench halubench , HaluEval li2023halueval , DefAn rahman2024defan and SimpleQA simpleQA are intended for factuality evaluation of LLMs, consisting of different types of questions such as reasoning, information retrieval and they vary in domains. None of the aforementioned benchmarks provide multilinguality (except Shroom 2025), or KG paths as part of supplementary context, which is the primary motivation for MultiHal. Furthermore GRAF cruaciun2024graf is a legal domain KG-based benchmark for Romanian language, although is limited by lack of multilinguality. MintakaQA sen-etal-2022-mintaka and MKQA mkqa datasets offers multilingual coverage as well as annotations of Wikidata entities for questions-answers sen-etal-2022-mintaka or only answers mkqa , but not full KG paths.
@@ 8 Limitations and Future Works
MultiHal is based around a multilingual question-answering task grounded with factual information; however ignoring use cases of multi-round dialogue and text summarization. Furthermore, our multilinguality can be considered limited in typological diversity ploeger-etal-2024-typological and bias towards European-centric languages. Multi-prompt evaluation has also been raised as an important component of LLM evaluation methodology whatSota ; maia2024efficient , which we do not cover, yet acknowledge as necessary, to test for LLM robustness and safety. Therefore, we propose expanding MultiHal’s downstream tasks and adding multiprompt evaluation and more diverse language coverage.
For evaluation of baseline experiments, we use three seperate models with no re-runs of random seeds. The evaluation of semantic similarity on a continuous scale makes the results hard to interpret across models, though still valid on a relative scale per model. Therefore, we propose future works of evaluating semantic similarity by framing it as a classification problem using a natural language inference sansford2024grapheval , or LLM-as-a-judge min-etal-2023-factscore .
@@ Appendix A MultiHal: Statistics and Schema
Appendix B Refusal Patterns
@@
Appendix C SPARQL Queries
@@
Listing 3: SPARQL query for 2-hop date retrieval subject-object. The \$OBJECT is a date string formatted as yyyy-mm-dd.
@@
Appendix D Numerical and Time-Based Properties
@@
Appendix E LLM Judge Prompts
@@
Appendix F Baseline Experiment Prompts
@@
Appendix G Overview of Preliminary Baseline Results per Domain and Dataset
@@
Appendix H Baseline Result Numerical Values
@@
Appendix I Example Problematic Data Points from Evaluation
@@
Appendix J CO2 Emission and Compute Resources Related to Experiments
@@
Estimations were conducted using the MachineLearning Impact calculator presented in [16].

### v2

@@ 2.4 Multilinguality
For enabling multilinguality for MultiHal, we employ the Nllb-200 3.3bn (Costa-Jussà et al., 2022) model and focus on its five well-performing European languages, namely German, Italian, French, Portuguese and Spanish. Our generation hyperparameters are specified in Table 2. Empirically, we found these hyperparameters to work the most optimal for our use case. We also noted that by separating the labels with semicolons yielded more accurate translations than having KG path labels purely whitespace separated, we attribute this to the improper grammatical structures that occur when label entities are not separated. We observe that Nllb-200’s output translations are generally of high quality, yet Nllb-200’s model does not always correctly output semicolon separation between the entities and predicates with respect to the English source.
@@ 3 Experimental Setup
The baseline experiments are set up using a prompt-based knowledge injection method. The prompt $\mathcal{P}$ is formatted as $\mathcal{P}=(\mathcal{K},\mathcal{Q})$, where $\mathcal{K}$ is knowledge in the form of a KG path and $\mathcal{Q}$ is the question of the data point, see Appendix G Listing 10 for KG-RAG and Listing 11 for vanilla QA for used prompts. We conduct experiments with and without knowledge $\mathcal{K}$ (KG-RAG and vanilla QA respectively) to measure the effectiveness of the factual information contained in the KG paths. We measure the semantic similarity between ground-truths and model predictions using Multilingual-MiniLM-L12-v2666https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 (Wang et al., 2024), the choice of the sentence embedding model is based on results in the MMTE benchmark (Enevoldsen et al., 2025), its multilingual capabilities, and comparatively small parameter count. Semantic similarity is computed by mean-pooling the last hidden states per token for each sentence, applying L2 normalization and computing the dot-product between the ground-truth and LLM prediction representations. For experimental conditions, we use Gemini 2.0 Flash, GPT-4o Mini, and Llama 3.3 70bn instruct models. Additionally we triangulate our semantic similarity results with NLI and hallucination detection. Inspired from Bang et al. (2025) we run hallucination detection on the English subsplit of MultiHal using HHEM-2.1 (Bao et al., 2024), and run NLI evaluation similar to Sansford et al. (2024); Zhang et al. (2024) based on a DeBERTa multilingual NLI model  [^7] https://huggingface.co/MoritzLaurer/mDeBERTa-v3-base-xnli-multilingual-nli-2mil7 namely due to model’s language coverage and its performance. The NLI and HHEM-2.1 models expect test pairs as hypothesis and premise. We form the premise as a prompt containing a question and ground-truth $\mathcal{P}=\{\mathcal{Q};\mathcal{G}\}$ and hypothesis is model response $\mathcal{A}$.
@@ 4.1 Preliminary Baselines
Considering that our methodology primarily relies on an LLM-judge for filtering and rating the quality of KG paths, we conduct a preliminary baseline test for the English subset to observe the performance of LLM judges. We conduct the experiment using a proportionally sampled subset of MultiHal, see Figure 2 for the data distribution. The goal of this is to gain insights on the expected quality of the KG paths.
@@
Metric Value
False Positives 11%
False Negatives 2.78%
IAA Cohen-Kappa 0.62
Table 4: GPT 4o-Mini overview of false positives, false negatives and IAA - all computed with respect to human judgement. Before computation all path scores are binarized as low (ratings 1, 2, 3) and high (ratings 4, 5) quality. Results are based on dataset subsplits in Figure 2.
From the results in Table 3 we see that paths judged by GPT 4o-Mini have a higher correlation with the semantic similarity. Given that low quality paths (rated 1-3) impair the LLM output quality and high quality paths (4-5) improve it (see Appendix M), we chose to run the full baseline experiments with GPT 4o-Mini. For a more in-depth breakdown of GPT 4o-Mini performance per dataset and domain, please refer to Appendix H. Table 4 showcases false positives and false negatives for GPT 4o-Mini as well as interannotator agreement with respect to human judgment as ground truth. IAA score between human and GPT 4o-Mini annotations, including a focus on misclassifications by GPT-4o which gives an indication of noise presence which is comparable to datasets and can be used for reproducibility.
@@ 4.2 Baseline Experiments
We report our baseline experiment results in Figure 3 for semantic similarity, Table 6 for aggregated NLI scores over languages and Table 6 for hallucination detection on the English subsplit, see Appendix O for a fine-grained overview of semantic similarity over domains and Appendix N for NLI results expanded over each language. The results showcase a consistent improvement of KG-RAG setting over QA for all evaluation metrics, indicating that the mined KG paths are meaningful for a model to generate a higher quality output. The result distributions in Figure 3 have statistically significant differences across all languages and all models between the QA and KG-RAG conditions, see Appendix L for more details. For a full numerical overview of Figure 3, see Appendix I.
@@
Model Task Ent Neut Contr
Gemini-2.0 Flash KG-RAG 68% 26% 6%
QA 52% 24% 24%
Llama-3.3-70b instr KG-RAG 71% 21% 8%
QA 45% 31% 23%
GPT-4o Mini KG-RAG 76% 16% 8%
QA 40% 32% 28%
Table 5: Aggregated NLI results over all MultiHal languages. Ent is entailment, Neut is neutral, Contr is contradicting
Model Task C H
Gemini-2.0 Flash KG-RAG 87% 13%
QA 58% 42%
Llama-3.3-70b instr KG-RAG 88% 12%
QA 55% 45%
GPT-4o Mini KG-RAG 89% 11%
QA 47% 53%
Table 6: Result overview with hallucination detection using HHEM-2.1. C (consistent), H (hallucinated)
@@ 5 Discussion
Overall, the results from Figure 3 and Tables 6, 6 depict a consistent improvement in a KG-RAG for all tested LLMs over QA; for detailed results of semantic similarity over domains refer to Appendix O. Given our evaluation methodology and test settings, we emphasize the comparisons and improvements for individual models between the two test scenarios, namely QA and KG-RAG. Therefore, we do not compare results across the models primarily due to varied parameter counts, and closed-source development of Gemini and GPT models. While scoping into specific domains in Table 15, we see that the performance fluctuates, although the foundational benchmarks SimpleQA, HaluEval, Defan and Shroom2024 contain approximately 95% of all data points for which we see consistent improvements on a per-model basis. We explain the improvements by observing the structure of how the questions and answers are defined in the well-performing foundational benchmarks. In a general case, as the Table 15 depicts the best-performing subsets, such as Defan, SimpleQA, HaluEval and Shroom2024, define the question explicitly and unambiguously with a single entity answer. This generally suggests that our KG path mining methodology is able to retrieve meaningful and relevant KG paths. Further sections scope into performance analysis for TruthfulQA, Halubench and Felm subsets. We supply some example problematic data points in Appendix J. Table 4 reports approximately 11% false positives which indicates levels of noise although this is still comparable and exceeds other QA datasets where upon analysis the noise ranges in 20-30% (Iqbal et al., 2024).
@@
Temporal, Leading, Suggestive and Reasoning Questions. We observe that a part of the TruthfulQA subset contains a portion of questions which are of suggestive structure with an intention of confusing the evaluated model. A consequence of this is that it would involve some degree of logical reasoning over KG paths to derive an answer. Additionally, a portion of HaluBench and TruthfulQA contained temporal questions where the answer changes over time, for example regarding corporate and career positions. Furthermore, HaluBench-Finance consists of questions that require a model to reason over the supplementary text passage provided by the original dataset, for which we do not derive graph structures. Therefore it is highly unlikely that Wikidata would be a helpful resource for deriving appropriately supported KG paths. Our evaluation pipeline could benefit from integration of a reasoning methodology similar as per Generate-on-Graph (Xu et al., 2024) or Think-on-Graph (Sun et al., 2023). Refer to Appendix J Listings 15 and 13 for explicit examples.
@@
Domains. Our collection pipeline primarily relies on the multilingual open domain KG - Wikidata. For domains such as HaluBench-Pubmed, TruthfulQA-Health and HaluBench-Covid, performance can be improved by utilizing medical domain knowledge graphs, for example PubMed (Xu et al., 2020) or PrimeKG (Chandak et al., 2023). We outline that the modularity of our pipeline allows for easy substitute of the KG endpoint.
@@
Sentence Embedding Limitations. We also note that our sentence embedding evaluation may not always accurately capture the semantics with respect to the question. In many cases, the ground-truth contained a repetition of the question, whereas our prompts contained instructions to answer concisely and explicitly, see Appendix G. Consequentially, some data points were evaluated with a relatively low semantic score even though the model responses directly, or with minimal deviations, answered the question. We note that TruthfulQA and Felm have been particularly affected by this as their ground-truth answers contain repetitions of the text but our model responses are more focused on single, explicit entities without linguistic overheads. Refer to Appendix J Listings 12 and 14. Therefore we triangulate our semantic similarity results with NLI and hallucination detection for increased confidence in our benchmark quality.
@@ 6 Related Work: Use of KGs in Datasets and Language Modeling
Multiple surveys discuss KG usage in the context of LLMs, particularly outlining future work roadmaps and synergy (Pan et al., 2023; 2024; Kau et al., 2024), discussing KGs in context of factuality, hallucination mitigation, multilinguality (Lavrinovics et al., 2025), and graph-retrieval augmented generation (Peng et al., 2024). We identify these as useful starting points for researchers new to the topic.
Language Modeling. Sansford et al. (2024); Rashad et al. (2024) make use of KG structures as part of their hallucination detection methodology by extracting graph structures from a given piece of text passage. Sun et al. (2023) approach involves reasoning over KGs and Srivastava et al. (2024) generates SPARQL queries from natural text. FactKG (Kim et al., 2023) and Fleek (Fatahi Bayat et al., 2023) propose methodologies using KGs to aid fact-checking. All the aforementioned language modeling approaches present KG information as in-context knowledge. However, in-context knowledge has limitations — particularly when there are conflicts between LLM’s internal knowledge and the provided context, or when there is limited transparency into how the model integrates and utilizes the external knowledge. An alternative approach is to encode the information as part of the model’s weights using adapter networks (Pfeiffer et al., 2023; Tian et al., 2024; Ribeiro et al., 2022).
Factually Oriented and KG-QA Datasets. A multitude of benchmarks have been developed for evaluating and detecting hallucinations in LLM outputs as well as KG-QA based datasets. Benchmarks such as Shroom2025 (Vázquez et al., 2025), Felm (Zhao et al., 2023), TruthfulQA (Lin et al., 2021), (Wei et al., 2024), HaluBench (Ravi et al., 2024), HaluEval (Li et al., 2023), DefAn (Rahman et al., 2024) and SimpleQA (Wei et al., 2024) are intended for factuality evaluation of LLMs, consisting of different types of questions such as reasoning, information retrieval and they vary in domains. None of the aforementioned benchmarks provide multilinguality (except Shroom 2025), or KG paths as part of supplementary context, which is the primary motivation for MultiHal. Furthermore GRAF (Crăciun et al., 2024) is a legal domain KG-based benchmark for Romanian language, although is limited by lack of multilinguality. MintakaQA (Sen et al., 2022) and MKQA (Longpre et al., 2021) datasets offers multilingual coverage as well as annotations of Wikidata entities for questions-answers (Sen et al., 2022) or only answers (Longpre et al., 2021), but not full KG paths.
@@ 8 Limitations and Future Works
MultiHal is based around a multilingual question-answering task grounded with factual information; however ignoring use cases of multi-round dialogue and text summarization. Furthermore, our multilinguality can be considered limited in typological diversity (Ploeger et al., 2024). We do not include a multi-prompt evaluation (Mizrahi et al., 2024; Maia Polo et al., 2024) and leave it for future expansion of this benchmark.
For evaluation of baseline experiments, we use three seperate models with no re-runs of random seeds. The evaluation of semantic similarity on a continuous scale makes the results hard to interpret across models, though still valid on a relative scale per model. None of our evaluation metrics provide a fine-grained overview pinpointing exact hallucinatory text spans.
@@
9 Reproducibility Statement
We outline that majority of our dataset collection pipeline uses interpretable methods such as SPARQL queries or open-source frameworks such as Falcon 2.0. Our automated path quality evaluation is based on LLM judge methodology using GPT-4o-Mini which is closed source, although we report interannotator agreement and false positives with respect to human judgement. In case of reproducibility, these numbers can be used as a guideline for evaluating reproduced collection with open sourced models. All of our evaluation metrics use open sourced models. We also open-source our data collection codebase which can be easily modified by replacing closed-source components for open-source ones and we diligently conduct a multi-faceted analysis which can be used as reference.
10 Ethics Statement
To the best of our knowledge our work does not raise any ethical concerns with respect to ICLR ethics policy. We comply with the dataset licenses and terms of use. To the best of our knowledge, the datasets do not contain personally identifiable information or sensitive content that could raise ethical concerns. See Appendix B for our LLM usage statement.
@@ Appendix A MultiHal: Statistics and Schema
Appendix B LLM Usage
In writing of this paper we use LLMs as a writing aid for formatting, grammar correction and paraphrasing. All the LLM generated writing is human cross-checked and cross-validated.
Appendix C Refusal Patterns
@@
Appendix D SPARQL Queries
@@
Listing 3: SPARQL query for 2-hop date retrieval subject-object. The \$OBJECT is a date string formatted as yyyy-$m$-$d$.
@@
Appendix E Numerical and Time-Based Properties
@@
Appendix F LLM Judge Prompts
@@
Appendix G Baseline Experiment Prompts
@@
Appendix H Overview of Preliminary Baseline Results per Domain and Dataset
@@
Appendix I Baseline Result Numerical Values
See Table 10 for numerical overview of our semantic similarity scores.
@@ Appendix I Baseline Result Numerical Values
Appendix J Example Problematic Data Points from Evaluation
@@
Appendix K CO2 Emission and Compute Resources Related to Experiments
@@
Estimations were conducted using the MachineLearning Impact calculator presented in (Lacoste et al., 2019).
Appendix L Statistical Significance Tests
We compute Shapiro-Wilk (SW) test for semantic score normality distribution and Cramer-von-Misses two-sample test for statistical significance between QA and KG-RAG distribution means of full MultiHal benchmark, see Table 12.
Language Model SW p-Value (KG-RAG) SW p-Value (QA) CvM p-Value
eng Gemini 2.0 1.53e-110 3.50e-76 2.54e-07
eng Llama 3.3 70B 2.48e-106 3.08e-71 3.40e-07
eng GPT 4o-Mini 1.02e-77 9.02e-71 1.47e-07
deu Gemini 2.0 3.23e-90 3.05e-73 1.01e-07
deu Llama 3.3 70B 1.13e-74 5.61e-69 1.07e-07
deu GPT 4o-Mini 7.94e-59 3.22e-63 8.26e-08
fra Gemini 2.0 4.58e-81 1.36e-69 1.39e-07
fra Llama 3.3 70B 6.84e-67 8.94e-70 1.28e-07
fra GPT 4o-Mini 5.01e-56 6.17e-68 6.68e-08
ita Gemini 2.0 4.84e-99 8.57e-75 2.34e-07
ita Llama 3.3 70B 3.09e-76 2.97e-65 1.24e-07
ita GPT 4o-Mini 3.38e-59 2.51e-62 1.17e-07
spa Gemini 2.0 2.02e-96 1.67e-72 1.38e-07
spa Llama 3.3 70B 1.94e-59 7.03e-64 9.78e-08
spa GPT 4o-Mini 5.34e-60 5.40e-63 7.37e-08
por Gemini 2.0 1.07e-104 3.20e-71 2.91e-07
por Llama 3.3 70B 3.65e-76 3.67e-69 1.61e-07
por GPT 4o-Mini 4.49e-65 5.20e-64 9.70e-08
Table 12: Shapiro-Wilk (SW) and Cramer-von Mises (CvM) by Model and Language for distribution of semantic scores.
Appendix M Path Quality Ablations
We compute further results on ablating the KG path quality levels from lowest quality (1) up to highest quality (5). Table 13 shows gradual shift towards high quality paths. The results are computed over the preliminary baseline distribution as per Figure 2.
Quality Levels Mean Std N
[1] 0.33 0.28 455
[1, 2] 0.37 0.31 1623
[1, 2, 3] 0.42 0.33 2046
[1, 2, 3, 4] 0.47 0.35 2451
[1, 2, 3, 4, 5] 0.51 0.36 2834
[2, 3, 4, 5] 0.55 0.36 2379
[3, 4, 5] 0.70 0.32 1211
[4, 5] 0.77 0.28 788
[5] 0.81 0.26 383
Table 13: Overview of semantic similarity with respect to KG path quality.
Appendix N NLI Test Results
Table 14 showcases NLI results expanded over languages with finer precission decimals with an aggregated total mean for a general overview.
Name Language Task Entailment Neutral Contradiction
google-gemini-2.0-flash-001 eng KG-RAG 65.70% 30.05% 4.25%
QA 56.13% 19.12% 24.76%
deu KG-RAG 67.79% 25.76% 6.45%
QA 49.96% 25.69% 24.35%
fra KG-RAG 58.32% 33.63% 8.06%
QA 44.95% 33.72% 21.33%
ita KG-RAG 75.53% 18.41% 6.06%
QA 53.79% 19.84% 26.37%
spa KG-RAG 71.86% 20.98% 7.16%
QA 54.05% 23.45% 22.50%
por KG-RAG 66.05% 28.28% 5.67%
QA 52.70% 23.55% 23.75%
meta-llama-llama-3.3-70b-instruct eng KG-RAG 68.91% 24.45% 6.64%
QA 48.50% 30.08% 21.42%
deu KG-RAG 71.31% 20.94% 7.76%
QA 44.33% 31.80% 23.88%
fra KG-RAG 64.70% 26.19% 9.11%
QA 42.08% 36.44% 21.47%
ita KG-RAG 74.36% 17.96% 7.68%
QA 45.63% 28.55% 25.82%
spa KG-RAG 71.01% 21.15% 7.84%
QA 43.61% 32.31% 24.08%
por KG-RAG 75.12% 18.09% 6.80%
QA 48.04% 29.12% 22.84%
openai-gpt-4o-mini eng KG-RAG 81.74% 12.10% 6.15%
QA 42.70% 26.33% 30.97%
deu KG-RAG 72.99% 18.42% 8.59%
QA 37.61% 35.84% 26.55%
fra KG-RAG 65.49% 24.48% 10.03%
QA 37.22% 39.77% 23.00%
ita KG-RAG 78.50% 13.65% 7.85%
QA 41.14% 28.42% 30.44%
spa KG-RAG 75.53% 15.74% 8.73%
QA 40.77% 32.31% 26.92%
por KG-RAG 80.93% 12.38% 6.69%
QA 41.44% 31.38% 27.18%
Total KG-RAG 71.44% 21.26% 7.31%
QA 45.81% 29.32% 24.87%
Table 14: NLI results over MultiHal benchmark.
Appendix O Fine-Grained Semantic Similarity Overview Across Domains
Table 15 showcases a breakdown of semantic similarity across domains of MultiHal.
Gemini 2.0 Flash Llama 3.3 70b instruct GPT 4o Mini
felm (wk) 73 17 0.63 0.794 -0.164 0.742 0.789 -0.047 0.853 0.826 0.027